Loc weight 14 pounds. The HCP prescribes an IV infusion of 5% dextrose in 1/4 normal saline to be infused at 8 ml / kg / HR. How many milliliters of the prescribed solution should the nurse Infuse each hour
Answer #1

weight of patient= 14 pounds

2.2 lb = 1 kg

14 lb = 14/2.2 kg = 6.36 kg

Rate of infusion= 8ml/kg/HR

Amount of prescribed solution the nurse should infuse each hour

= 6.36 × 8 = 50.88 ml